# Editable-DeepSC: Reliable Cross-Modal Semantic Communications for Facial Editing

### v3

@@ Abstract
Real-time computer vision (CV) plays a crucial role in various real-world applications, whose performance is highly dependent on communication networks. Nonetheless, the data-oriented characteristics of conventional communications often do not align with the special needs of real-time CV tasks. To alleviate this issue, the recently emerged semantic communications only transmit task-related semantic information and exhibit a promising landscape to address this problem. However, the communication challenges associated with Semantic Facial Editing, one of the most important real-time CV applications on social media, still remain largely unexplored. In this paper, we fill this gap by proposing Editable-DeepSC, a novel cross-modal semantic communication approach for facial editing. Firstly, we theoretically discuss different transmission schemes that separately handle communications and editings, and emphasize the necessity of Joint Editing-Channel Coding (JECC) via iterative attributes matching, which integrates editings into the communication chain to preserve more semantic mutual information. To compactly represent the high-dimensional data, we leverage inversion methods via pre-trained StyleGAN priors for semantic coding. To tackle the dynamic channel noise conditions, we propose SNR-aware channel coding via model fine-tuning. Extensive experiments indicate that Editable-DeepSC can achieve superior editings while significantly saving the transmission bandwidth, even under high-resolution and out-of-distribution (OOD) settings.
@@
semantic communications, cross-modal data, facial editing tasks, generative adversarial networks, deep learning
@@ 1 Introduction
In the era of rapid technological advancement, real-time computer vision (CV) plays a pivotal role in various real-world applications, such as digital manufacturing [1], telemedicine [2], robotics [3], and metaverse [4]. The performance of real-time CV tasks is heavily dependent on the underlying communication networks. These traditional communication systems are primarily designed for data-oriented purpose and focus on metrics that measure the data transmission instead of the tasks execution performance. In other words, the original data is expected to be completely recovered at the receiver side regardless of the users’ usage about it. Nonetheless, this data-oriented design principle is not always suitable for the specific needs of real-time CV tasks, as different tasks at the receiver side would require different types of extracted semantic information from the original data, which should be encoded and transmitted with different importance. Reconstructing every part of the original data equally will undoubtedly waste the limited transmission bandwidth.
To alleviate this issue, the recently emerged semantic communications [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] (also known as task-oriented communications) only transmit the specific semantic information suitable for real-time CV tasks and exhibit a promising landscape to address this problem. By bridging the gap between bit-level transmission and task-oriented requirements, semantic communications can exploit much more of the scarce bandwidth to transmit task-related information, thus realizing better real-time CV tasks execution performance. On the other hand, Semantic Facial Editing [17, 18, 19, 20, 21, 22], which stands out as one of the most important real-time CV tasks for user interaction and personalization, encounters various communication challenges when promoting its real-world deployment. As shown in Figure 1, in many famous social platforms (e.g., Facebook, Instagram), before transmitting their own photos to the remote servers, users may wish to flexibly edit the original data according to their personal needs, such as adding smile to a face or altering the transparency of eyeglasses. Furthermore, such personal requirements can be fulfilled by providing dialogues to enable a more conversational and interactive experience. A crucial question still remains largely unexplored in existing works: How to better transmit the facial semantics while also dynamically editing them according to the users’ needs?
@@ 3. $Q_{3}$:
    How to better tackle the dynamically varying channel noise conditions, so that the model can generalize well to different circumstances of communication capabilities?
To address the above challenges, we first theoretically discuss different transmission schemes that separately handle communications and editings, where both the cases of editings conducted at the receiver side (i.e., editings after the channel transfer) and editings conducted at the transmitter side (i.e., editings before the channel transfer) are taken into consideration. We find that by separately dealing with communications and editings, these schemes actually introduce more encoding and decoding steps, increase the data processing procedures, and are more likely to lose semantic mutual information, which will undermine the communication efficiency and eventually reduce the quality of edited images at the receiver side. In light of this, we emphasize the necessity to integrate editings into the communication chain, jointly optimize these two aspects, and minimize the data processing procedures to preserve more semantic mutual information.
Then, we propose semantic coding based on Generative Adversarial Networks (GAN) inversion methods [23, 24]. By leveraging the abundant generative priors within the pre-trained StyleGAN [25], we encode the input images into the GAN latent space and map them to low-dimensional representations. Thus, we compactly represent the high-dimensional input data and considerably decrease the Channel Bandwidth Ratio (CBR) required to transmit the edited facial semantics. This addresses the aforementioned $Q_{1}$.
Inspired by the fact that DeepJSCC [26] combines source coding and channel coding to jointly optimize data compression and error correction, we propose Joint Editing-Channel Coding (JECC) based on iterative attributes matching. We design the JECC codecs by Fully Connected Layers, which iteratively exert minor modifications on the previous results of semantic coding to realize semantic editing and channel coding at the same time. Specifically, the current degrees of image attributes output by the pre-trained attribute predictor are compared with expected degrees after each iteration of modification to ensure fine-grained editings, where only the targeted attributes (e.g., bangs, eyeglasses) are edited without influencing other untargeted attributes. Moreover, the JECC codecs are trained through noisy channels to guarantee their basic robustness against channel noises in addition to the semantic editing abilities. Thus, we preserve more semantic mutual information due to the reduced data processing procedures of JECC and precisely perform fine-grained editings only on the targeted semantics. This addresses the aforementioned $Q_{2}$.
Benefiting from the recent successful applications of transfer learning techniques [27, 28], we propose SNR-aware channel coding based on model fine-tuning. Specifically, we add two lightweight trainable adapters that do not change the shapes of inputs and outputs within the JECC codecs. When fine-tuning the models, only the parameters of these two adapters are adjusted to capture the distribution of new noise conditions. Meanwhile, the rest of the parameters are frozen to avoid forgetting the previously learned priors because of the fine-tuning. Thus, we efficiently adapt the model to the varying circumstances of communication capabilities by fine-tuning only a very small portion of the total parameters. This addresses the aforementioned $Q_{3}$.
@@
To verify the effectiveness of our method, we conduct extensive experiments under various levels of SNRs and compare Editable-DeepSC with the baselines that separately handle communications and editings, where both editings after the channel transfer and before the channel transfer are evaluated. We also consider more rigorous and realistic simulation settings where the transmitter side users may hold high-resolution images or out-of-distribution (OOD) images. Numerous results indicate that Editable-DeepSC can achieve superior editing effects compared to the baseline methods while significantly saving the transmission bandwidth. For instance, in high-resolution settings, Editable-DeepSC only consumes less than $2\%$ of the baseline methods’ CBR, but still achieves the best editing performance both quantitatively and qualitatively. Furthermore, in the fine-tuning stage, Editable-DeepSC only adjusts $2.65\%$ of the total parameters, but still considerably improves the editing effects under low SNRs and protects the edited semantics from channel noise corruptions.
@@
  • (1)
    We theoretically analyze different transmission schemes that separately handle communications and editings. We discover that they actually increase the data processing procedures and are more likely to lose semantic mutual information. In light of this, we emphasize the necessity to integrate editings into the communication chain, jointly optimize these two aspects, and preserve more semantic mutual information.
  • (2)
    We propose semantic coding via GAN inversion. By leveraging the abundant generative priors within the pre-trained StyleGAN, we compactly compress the high-dimensional input images to low-dimensional representations to better reduce the bandwidth consumption.
  • (3)
    We propose Joint Editing-Channel Coding (JECC) via iterative attributes matching. We iteratively exert minor modifications on the results of semantic coding to reduce the data processing procedures for semantic mutual information preservation and precisely perform fine-grained editings only on the targeted semantics.
  • (4)
    We propose SNR-aware channel coding via model fine-tuning. We introduce two lightweight trainable adapters to capture the distribution of new noise conditions and freeze the rest parameters to avoid forgetting the priors, which helps the model generalize well to the varying circumstances of communication capabilities.
  • (5)
    Extensive experimental results demonstrate that Editable-DeepSC can achieve superior performance compared to existing methods in terms of editing effects and transmission efficiency, even under more practical settings with high-resolution and out-of-distribution (OOD) images.
@@ 2.1 Semantic Facial Editing
Semantic Facial Editing aims to provide personalized facial manipulations specified by the users, and has emerged as one of the most important real-time CV tasks for its wide range of applications, including virtual content creation [29], augmented reality (AR) [30], and online conferencing [31]. This technology enables users to make intuitive edits to facial attributes such as expression, age, hairstyle, offering both creative flexibility and enhanced user experiences.
In the early stage of this research field’s development, many methods [17, 18, 19] focus on editing specific attributes, and rely heavily on extra manually provided knowledge such as landmarks. Recently, latent space based facial editing methods [20, 21, 22] are proposed due to the advancement of generative models like StyleGAN [25]. These methods typically strive to find semantically meaningful directions in the latent space of pre-trained generative models, so that shifting along the latent space would be able to achieve the desired editing in the vision space. InterFaceGAN [20] finds a hyperplane in the latent space to disentangle the facial semantics, and uses the normal vector of the hyperplane as the editing direction. Zhuang et al. [21] proposed to learn a mapping network to generate identity-specific directions within the latent space. However, both the two methods [20, 21] neglect language guidance and lack conversational interaction. To alleviate this issue, Jiang et al. [22] proposed the SOTA approach Talk2Edit, which enables customized editings via human language feedback.
@@ 2.2 Task-Oriented Semantic Communications
Semantic communications prioritize the transmission of task-related semantic information rather than the original data with enhanced communication efficiency, and have showed great potential across various real-time CV tasks.
Single-modal real-time image tasks. Zhang et al. [5] proposed a DNN-based image semantic communication system, which only transmits the gradients back to the transmitter during training to protect the receiver’s privacy of downstream tasks. From the perspective of Reinforcement Learning (RL) [32], Huang et al. [6] adopted an adaptive semantic coding approach with an RL-based bit allocation model for better semantic similarity and image perception. Dong et al. [7] proposed the concept of semantic slice models, which can flexibly adapt the models to different requirements of model performance, channel situations and transmission goals.
Single-modal real-time video tasks. Tung et al. [8] first proposed to directly map video signals to channel symbols, and utilized RL [32] to optimize the allocation of bandwidth among video frames. Jiang et al. [9] proposed a video semantic coding network based on keypoints transmission, and considerably reduced resources consumption without losing the main details. Wang et al. [10] exploited nonlinear transform and conditional coding architecture to adaptively extract semantic features across video frames, and surpassed traditional wireless video coded transmission schemes.
Multi-modal real-time CV tasks. Xie et al. [11] proposed a multi-modal semantic communication system for Visual Question Answering (VQA) tasks, which utilizes MAC networks [33] to deal with the interconnected cross-modal data and generate the answers to VQA tasks. Xie et al. [12] further considered the multi-modal multi-user transmission scenarios and proposed a unified framework to effectively reduce the mutual interference between different users. Wang et al. [13] utilized Transformer [34] to design a distributed Audio-Visual Parsing (AVP) network under transmission scenarios with superior multi-modal parsing performance.
@@ 3 System Model
Figure 2: Overview of the proposed framework. Editable-DeepSC mainly consists of the Text Semantic Encoder, the Image Semantic Encoder, the Joint Editing-Channel Encoder, and the Joint Semantic-Channel Decoder, where channel noise corruptions from the real world are also taken into consideration.
As shown in Figure 2, our proposed Editable-DeepSC mainly consists of the Text Semantic Encoder, the Image Semantic Encoder, the Joint Editing-Channel Encoder, and the Joint Semantic-Channel Decoder, where channel noise corruptions from the real world are also taken into consideration.

### v4

@@ Abstract
Interactive computer vision (CV) plays a crucial role in various real-world applications, whose performance is highly dependent on communication networks. Nonetheless, the data-oriented characteristics of conventional communications often do not align with the special needs of interactive CV tasks. To alleviate this issue, the recently emerged semantic communications only transmit task-related semantic information and exhibit a promising landscape to address this problem. However, the communication challenges associated with Semantic Facial Editing, one of the most important interactive CV applications on social media, still remain largely unexplored. In this paper, we fill this gap by proposing Editable-DeepSC, a novel cross-modal semantic communication approach for facial editing. Firstly, we theoretically discuss different transmission schemes that separately handle communications and editings, and emphasize the necessity of Joint Editing-Channel Coding (JECC) via iterative attributes matching, which integrates editings into the communication chain to preserve more semantic mutual information. To compactly represent the high-dimensional data, we leverage inversion methods via pre-trained StyleGAN priors for semantic coding. To tackle the dynamic channel noise conditions, we propose SNR-aware channel coding via model fine-tuning. Extensive experiments indicate that Editable-DeepSC can achieve superior editings while significantly saving the transmission bandwidth, even under high-resolution and out-of-distribution (OOD) settings.
@@
Semantic Communications, Cross-Modal Data, Facial Editing Tasks, Generative Adversarial Networks, Deep Learning.
@@ 1 Introduction
In the era of rapid technological advancement, interactive computer vision (CV) plays a pivotal role in various real-world applications, such as digital manufacturing [8], telemedicine [1], robotics [39], and metaverse [44]. The performance of interactive CV tasks is heavily dependent on the underlying communication networks. These traditional communication systems are primarily designed for data-oriented purpose and focus on metrics that measure the data transmission instead of tasks execution performance. In other words, the original data is expected to be completely recovered at the receiver side regardless of the users’ usage about it. Nonetheless, this data-oriented design principle is not always suitable for the specific needs of interactive CV tasks, as different tasks at the receiver side would require different types of extracted semantic information from the original data, which should be encoded and transmitted with different importance. Reconstructing every part of the original data equally will undoubtedly waste the limited transmission bandwidth.
To alleviate this issue, the recently emerged semantic communications [58, 22, 47, 52, 46, 56] (also known as task-oriented communications) only transmit the specific semantic information suitable for interactive CV tasks and exhibit a promising landscape to address this problem. By bridging the gap between bit-level transmission and task-oriented requirements, semantic communications can exploit much more of the scarce bandwidth to transmit task-related information, thus realizing better interactive CV tasks execution performance. On the other hand, Semantic Facial Editing [54, 31, 48, 38, 60, 24], which stands out as one of the most important interactive CV tasks for user interaction and personalization, encounters various communication challenges when promoting its real-world deployment. As shown in Figure 1, in many famous social platforms (e.g., Facebook, Instagram), before transmitting their own photos to the remote servers, users may wish to flexibly edit the original data according to their personal needs, such as adding smile to a face or altering the transparency of eyeglasses. Furthermore, such personal requirements can be fulfilled by providing dialogues to enable a more conversational and interactive experience. A crucial question still remains largely unexplored in existing works: How to better transmit the facial semantics while also dynamically editing them according to the users’ needs?
@@ 3. $Q_{3}$:
    How to better tackle the dynamically varying channel conditions, so that the model can generalize well to different circumstances of communication capabilities?
To address the above challenges, we propose Editable-DeepSC, a novel framework that integrates facial editing directly into the semantic communication chain. Unlike conventional separate schemes (where editing is performed either before or after transmission) that introduce redundant data processing procedures and risk losing semantic mutual information, our approach employs a joint optimization strategy to preserve semantic fidelity and enhance editing quality.
Editable-DeepSC operates through three cohesive stages. First, it leverages generative priors for efficient semantic coding, dramatically compressing high-dimensional facial images into low-dimensional representations to tackle $Q_{1}$. Second, it introduces a Joint Editing-Channel Coding (JECC) mechanism that simultaneously performs semantic editing and ensures robustness against channel noise, addressing the fine-grained control requirements of $Q_{2}$. Finally, an SNR-aware adaptation mechanism enables the model to efficiently generalize across diverse channel conditions by fine-tuning only a minimal set of parameters, effectively solving $Q_{3}$.
@@
To verify the effectiveness of Editable-DeepSC, we conduct extensive experiments under various levels of SNRs and compare it with the baselines that separately handle communications and editings, where both editings after the channel transfer and before the channel transfer are evaluated. We also consider more rigorous and realistic simulation settings where the transmitter side users may hold high-resolution images or out-of-distribution (OOD) images. Numerous results indicate that Editable-DeepSC achieves superior editing effects compared to the baselines while significantly saving the transmission bandwidth. For instance, in high-resolution settings, Editable-DeepSC only consumes less than $2\%$ of the baseline methods’ bandwidth, but still achieves the best editing performance both quantitatively and qualitatively. Furthermore, in the fine-tuning stage, Editable-DeepSC only adjusts $2.65\%$ of the total parameters, but still considerably improves the editing effects under low SNRs and protects the edited semantics from channel noise corruptions.
@@
  • We theoretically analyze different transmission schemes that separately handle communications and editings, encompassing both transmitter-side (pre-channel) and receiver-side (post-channel) editings. We discover that they actually increase the data processing procedures and are more likely to lose semantic mutual information, undermining efficiency and ultimately degrading received image quality. In light of this, we emphasize the necessity to integrate editings into the communication chain, jointly optimize these two aspects, and minimize the data processing procedures to preserve more semantic mutual information.
  • We propose semantic coding via Generative Adversarial Networks (GAN) inversion. By leveraging the abundant generative priors within the pre-trained StyleGAN, we encode the high-dimensional input images into the GAN latent space and map them to low-dimensional representations. Thus, we compactly represent the high-dimensional input data and considerably decrease the Channel Bandwidth Ratio (CBR) required to transmit the edited facial semantics.
  • We introduce Joint Editing-Channel Coding (JECC) via iterative attributes matching. By employing low-complexity Fully Connected Layers and an attribute-matching loop, we iteratively exert minor modifications on the results of semantic coding to realize semantic editing and channel coding at the same time. Thus, we reduce the data processing procedures for semantic mutual information preservation and precisely perform fine-grained editings only on the targeted semantics, without influencing other untargeted ones.
  • We present SNR-aware channel coding via model fine-tuning. We introduce two lightweight trainable adapters (which do not change the shapes of inputs and outputs within the JECC codecs) to capture the distribution of new noise conditions. When fine-tuning the models, only the parameters of these two adapters are adjusted to learn the characteristics of new transmission environments. Meanwhile, the rest of the parameters are frozen to avoid forgetting the previously learned priors due to the fine-tuning. This helps the model generalize well to the varying communication capabilities via minimal parameters fine-tuning.
  • Extensive experiments demonstrate that Editable-DeepSC achieves superior performance compared to existing methods on editing effects and transmission efficiency, even under more practical settings with high-resolution and out-of-distribution (OOD) images.
@@ 2.1 Semantic Facial Editing
Semantic Facial Editing aims to provide personalized facial manipulations specified by the users, and has emerged as one of the most important interactive CV tasks for its wide range of applications, including virtual content creation [45], augmented reality (AR) [2], and online conferencing [35]. This technology enables users to make intuitive edits to facial attributes such as expression, age, hairstyle, offering both creative flexibility and enhanced user experiences.
In the early stage of this area, many methods [31, 48] focus on editing specific attributes, and rely heavily on extra manually provided knowledge such as landmarks. Recently, latent space based facial editing methods [38, 60, 24] are proposed due to the advancement of generative models like StyleGAN [27]. These methods typically strive to find semantically meaningful directions in the latent space of pre-trained generative models, so that shifting along the latent space would be able to achieve the desired editing in the vision space. InterFaceGAN [38] finds a hyperplane in the latent space to disentangle the facial semantics, and uses the normal vector of the hyperplane as the editing direction. Zhuang et al. [60] proposed to learn a mapping network to generate identity-specific directions within the latent space. However, both the two methods [38, 60] neglect language guidance and lack conversational interaction. To alleviate this issue, Jiang et al. [24] proposed Talk2Edit, which enables customized editings via human language feedback.
While diffusion models have recently gained significant popularity in image editing tasks [28], and have indeed been explored in semantic communications [50], we specifically choose the GAN architecture for Editable-DeepSC due to several practical advantages it offers. First, GAN-based models provide a well-structured latent space that enables more precise and disentangled control over facial attributes [27], which is crucial for achieving the fine-grained editing capabilities required by our framework. Second, the iterative denoising process of diffusion models typically requires multiple forward passes during inference [28], which would introduce substantial latency incompatible with interactive scenarios. In contrast, GAN models can be optimized to achieve a better balance between efficiency and quality, making it more suitable for our communication-efficient framework.
@@ 2.2 Task-Oriented Semantic Communications
Semantic communications prioritize the transmission of task-related semantic information rather than the original data with enhanced communication efficiency, and have showed great potential across various CV tasks.
Single-modal semantic communications for image tasks. Zhang et al. [58] proposed a DNN-based image semantic communication system, which only transmits the gradients back to the transmitter during training to protect the receiver’s privacy of downstream tasks. Xie et al. [53] proposed an image semantic communication method with the joint control of shallow and deep networks. Han et al. [17] proposed Standards-Compatible Semantic Communication (SCSC), an end-to-end image transmission approach with learnable preprocessing-empowered and precoder/combiner-enhanced modules for MIMO channels.
Single-modal semantic communications for video tasks. Jiang et al. [22] proposed a video semantic coding network based on keypoints transmission, and considerably reduced resources consumption without losing the main details. Wang et al. [47] exploited nonlinear transform and conditional coding architecture to adaptively extract semantic features across video frames, and surpassed traditional wireless video coded transmission schemes. Li et al. [29] proposed a video semantic communication system based on major object extraction and contextual video encoding, which can achieve efficient video transmission for massive communication.
Semantic communications for multi-modal/cross-modal CV tasks. Xie et al. [52] proposed a multi-modal semantic communication system for Visual Question Answering (VQA) tasks, which utilizes MAC networks [20] to deal with the interconnected cross-modal data and generate the answers to VQA tasks. Jiang et al. [21] proposed a novel Vision-Language Model-based Cross-modal Semantic Communication (VLM-CSC) system, which consists of three innovative components: Cross-modal Knowledge Base (CKB), Memory-assisted Encoder and Decoder (MED), and Noise Attention Module (NAM). VLM-CSC effectively addresses the challenges in dynamic environments such as low information density and catastrophic forgetting. Tian et al. [40] proposed a synchronous multi-modal semantic communication System with packet-level coding (SyncSC) for facial video and speech transmission, which reduces transmission overhead while ensuring high-quality synchronous transmission of video and speech over the packet loss network.
@@ 3 System Model
Table 1: Definitions of used symbols throughout this paper.
Symbol Description Symbol Description $S_{I}$ Input image, $\mathbb{R}^{1\times C\times W\times H}$ $S_{T}$ Input sentence, $\mathbb{R}^{1\times L_{2}\times L_{3}}$ $\hat{S}_{I^{\prime}}$ Reconstructed (edited) image, $\mathbb{R}^{1\times C\times W\times H}$ $E_{I}$ Image features, $\mathbb{R}^{1\times L_{1}}$ $E_{T}$ Text features, $\mathbb{R}^{1\times L_{4}}$ $X$ Transmitted signal, $\mathbb{C}^{1\times k}$ $Y$ Received signal, $\mathbb{C}^{1\times k}$ $X^{\prime},Y^{\prime}$ Signal after/before adapter, $\mathbb{C}^{1\times k}$ $h$ Channel coefficient, $\mathbb{C}$ NN Gaussian noise, $\mathbb{C}^{1\times k}$ $\mathcal{SE_{I}}(\cdot;\alpha)$ Image Semantic Encoder $\mathcal{SE_{T}}(\cdot;\beta)$ Text Semantic Encoder $\mathcal{JECE}(\cdot,\cdot;\gamma)$ Joint Editing-Channel Encoder $\mathcal{JSCD}(\cdot;\theta)$ Joint Semantic-Channel Decoder $\mathcal{A}_{CE}(\cdot;\phi)$ Channel Encoder Adapter $\mathcal{A}_{CD}(\cdot;\psi)$ Channel Decoder Adapter $G(\cdot)$ Pre-trained StyleGAN Generator $\mathcal{M}(\cdot;\gamma)$ Fully Connected Layers in $\mathcal{JECE}(\cdot,\cdot;\gamma)$ $P(\cdot)$ Pre-trained attribute predictor $D(\cdot)$ StyleGAN Discriminator $F(\cdot)$ Pre-trained Face recognition model $z$ Latent vector in StyleGAN space $z^{(t)}$ Latent after $t$-th inversion iteration $E_{I^{\prime}}^{(t)}$ Latent after $t$-th editing iteration $I_{g}$ Image generated by $G(z)$ $J(z)$ Distortion in inversion $\lambda_{inv}$ Weight for pixel-level distortion $\eta_{inv}$ Learning rate for inversion $r_{inv}$ Total inversion iterations $a_{i}$ Degree of the $i$-th attribute $m$ Total number of attributes $u$ Maximum degree value for each attribute $i_{tar}$ Index of targeted attribute $d_{tar}$ Editing direction (1-1 or +1+1) $\Delta_{tar}$ Degree of modification $a_{exp}$ Expected attribute degree $r_{edit}$ Maximum editing iterations $b_{ij}$ One-hot of expected attribute degrees $p_{ij}$ Predicted probability of attribute $i$ degree $j$ $\mathcal{L}_{pred}$ Predictor loss $\mathcal{L}_{id}$ Identity preservation loss $\mathcal{L}_{disc}$ Discriminator loss $\lambda_{pred},\lambda_{id},\lambda_{disc}$ Loss weights for training JECC codecs $I_{a},I_{b}$ Images after/before editing $n_{ft}$ Number of fine-tuning images $\mathcal{L}_{ft}$ Fine-tuning loss $\eta_{ft}$ Fine-tuning learning rate $r_{ft}$ Fine-tuning iterations $P_{max}$ Average transmission power constraint $\rho$ Channel Bandwidth Ratio (CBR) $K_{0}$ Distribution of original images $K_{n}$ Distribution of final edited images $K_{s}$ Distribution of intermediate representations $\mathcal{I}(\cdot;\cdot)$ Mutual information
As shown in Figure 2, our proposed Editable-DeepSC mainly consists of the Text Semantic Encoder, the Image Semantic Encoder, the Joint Editing-Channel Encoder, and the Joint Semantic-Channel Decoder, where channel noise corruptions from the real world are also taken into consideration. A summary of used mathematical symbols throughout this paper is shown in Table 1.